Ace rents a compact car at a daily rate of 40.95 plus.47 per mile. Deuce rents a compact car for 43.95 plus.42 per mile. For what mileage is the cost the same?
thanks

Answer by stanbon(75887)

Ace rents a compact car at a daily rate of 40.95 plus.47 per mile. Deuce rents a compact car for 43.95 plus.42 per mile. For what mileage is the cost the same?
----
Ace cost: A(x) = 40.95 + 0.47x
---
Deuce cost: D(x) = 43.95 + 0.42x
----
Solve A(x) = D(x) for "x":
40.95 + 0.47x = 43.95 + 0.42x
---
0.05x = 3
x = 60 miles (the cost is the same if the car is drawn 60 miles)
==============
Cheers,
Stan H.